News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 17 March 2017 9.00am EST

During the last week the Chandra observing schedule was replanned following the transition to Bright Star Hold on Mar 9 which was discussed in the status report last week. Loads for a replanned schedule were uplinked on Mar 10 with 82.7 ks of science loss. Observations of The Coma Cluster and Markarian 421 were impacted by the interruption and will be rescheduled.

A second replan of the observing schedule was done to reflect a new planning guideline that all Normal Point Mode dwells must be at least 20 minutes in duration. This is in response to continuing occurrence of gyro bias variations similar to those suspected in the two recent safing events. The replan had no impact on science.

A Chandra press release was issued on Mar 13 concerning X9 in 47 Tucanae. New Chandra data support the hypothesis that X9 consists of a black hole pulling gas off of a white dwarf companion. The X-rays change periodically over about 28 minutes and show evidence for large amounts of oxygen in the system. For details see:

The Chandra proposal deadline passed on Mar 16, it was postponed by one day due to a snowstorm on the East Coast of the U.S.

The schedule of targets for the next week is shown below and includes an observation of Jupiter coordinated with Juno, an observation of HD 159532 coordinated with HST, and an observation of GRS 1915+105 coordinated with NuSTAR, ATCA and the VLT.

        SNR 0509-68.7          ACIS-S       Mar 20 
        PS12zc                 ACIS-S       Mar 21
        IC 4040                ACIS-S   
        Cygnus A - W Hotspot   ACIS-I   
        SDSSJ134701.54+21540   ACIS-S   
        SDSSJ1030+0524         ACIS-I   
        Radiation Belt                      Mar 22   
        E0102-72               ACIS-S       Mar 23
        HD159532               HRC-I    
        W0533-3401             ACIS-S   
        2MASS J03103891+2540   ACIS-S   
        SDSSJ1030+0524         ACIS-I       Mar 24
        SNR 0509-68.7          ACIS-S   
        Cygnus A - W Hotspot   ACIS-I       Mar 25
        Radiation Belt
        SDSSJ1030+0524         ACIS-I       Mar 26
        2MASS J041146+132416   ACIS-S  
        E0102-72               ACIS-S    
        Cygnus A - W Hotspot   ACIS-I  
        GRS 1915+105           ACIS-S/HETG  Mar 27
        JUPITER                HRC-I   


All spacecraft subsystems continued to support nominal operations.

Return to Status Reports